A rare isotope of a nuclear material is very unstable, decaying at a rate of 12% each second. Find how much isotope remains 10 seconds after 4 grams of the isotope is created. Use y=4(2.7)^0.12t, where t represents time in seconds..
There would be how many grams of the isotope.
(Round to the nearest tenth as needed)

Answer by nerdybill(7384)

A rare isotope of a nuclear material is very unstable, decaying at a rate of 12% each second. Find how much isotope remains 10 seconds after 4 grams of the isotope is created. Use y=4(2.7)^0.12t, where t represents time in seconds..
.
In the given equation: (I think you need a negative sign)
y=4(2.7)^-0.12t
y is the amount (grams) left
after t (seconds)
.
simply replace t with 10 and equate:
y=4(2.7)^-0.12t
y=4(2.7)^-0.12(10)
y=4(2.7)^-1.2
y=4(0.30364)
y=1.2 grams